OpenStudy (tiffany_rhodes):

First make a substitution and then use IBP to evaluate the integral -4t^3*cos(t^2)dt from the sqr(pi/2) to the sqr (pi)?

OpenStudy (tiffany_rhodes):

I let u=t^2 and du=2tdt and dt=du/2t.

OpenStudy (anonymous):

use t^2 as z

OpenStudy (anonymous):

yup going in correct way

OpenStudy (anonymous):

integral of -4t^3*cos(t^2)dt if u=t^2du=2tdt then question can be written as -(2t^2)(2t) cos(t^2)dt the question transforms to -2ucosudu and the n use integration by parts

OpenStudy (tiffany_rhodes):

Okay, thank you!